# SOUL: Unlocking the Power of Second-Order Optimization for LLM Unlearning

### v1

@@ 5.1 Experiment setups
As for model configurations, We use the OPT-1.3B Zhang et al. (2022b) and LLaMA2-7b-chat Touvron et al. (2023) as our base models. In the TOFU task, we specifically utilized the fine-tuned LLaMA2-7b-chat model. In the copyright removal task, we fine-tuned both the OPT-1.3B and LLaMA2-7b-chat models on the Harry Potter book series dataset Eldan and Russinovich (2023) to simulate the generation of copyrighted information. In the detoxification task, the original models were used without any additional modification. Further information regarding model preparation and configuration can be found in Appendix A.1.
@@ 5.3 Results on copyright removal
Table 3 presents the unlearning efficacy and model utility of the proposed SO unlearning methods and baselines in the task of ‘Who’s Harry Potter’ copyrighted information removal across two LLMs fine-tuned on the Harry Potter book series dataset Eldan and Russinovich (2023). Consistent with our observations in the TOFU task, SOUL substantially improves the unlearning efficacy. For example, the comparison between FO-GradDiff and SO-GradDiff shows a notable decrease in BLEU score (by 0.21) at a prompt length of 300 in the LLaMA2-7B-chat model. This decrease suggests that the generated texts deviate further from the original book’s content. Furthermore, the enhancements observed in both perplexity (PPL) and zero-shot accuracy with SOUL over FO unlearning highlight a superior balance between forget efficacy and utility preservation. Similar to the TOFU task, the GA method struggles to balance forget efficacy with utility preservation. Despite achieving the lowest scores on the LLaMA2-7B-chat model, it results in notably poor utility, as evidenced by a perplexity of 15.66, substantially higher than other methods. Table A5 in Appendix A.7 showcases visualization examples, further demonstrating the enhanced performance of SOUL.
@@ 5.4 Results on LLM detoxification
In Table 4, we demonstrate that the proposed SO unlearning methods effectively reduce the toxicity score on both the Real Toxicity Prompts and PKU-SafeRLHF datasets while maintaining or even improving utility. For instance, in the LLaMA2-7B-chat model, SO-PO achieved a clear reduction in the toxic score on the PKU-SafeRLHF dataset and showed enhanced performance in zero-shot accuracy compared to FO-PO. This indicates improved unlearning efficacy of SOUL without sacrificing model utility. In addition, Table A6 includes visualizations that exemplify the outputs after the application of unlearning to the LLaMA2-7B-chat models. These visualizations further corroborate that SO optimizers improve unlearning efficacy, particularly highlighting that SO-PO achieves the most effective unlearning performance.
@@ 7 Limitations
Model scale limitation: Our experiments primarily focused on models like OPT-1.3B and LLaMA2-7b-chat. However, larger models, such as expanded variants of LLaMA, are increasingly common. The computational demands and unique characteristics of these larger models may affect the applicability or effectiveness of second-order unlearning techniques. Further investigation on larger-scale models is warranted to understand their behavior under second-order optimization.
@@ Appendix A Additional Experimental Details and Results
We selected the OPT-1.3B Zhang et al. (2022a) and LLaMA2-7b-chat Touvron et al. (2023) as foundational models for our study. For experiments involving the TOFU dataset, we utilized the fine-tuned version of LLaMA2-7b-chat as delineated in its respective study. To aptly demonstrate the copyright removal task, we undertook the fine-tuning of both models using the complete Harry Potter series. The fine-tuning procedure for the OPT-1.3B model involved a learning rate of $5\times 10^{-5}$ and a batch size of 2. Conversely, for LLaMA2-7b-chat, we applied Low-Rank Adaptation (LoRA) fine-tuning with a learning rate of $1\times 10^{-4}$ and the same batch size. AdamW served as the optimizer for preparing these models. For the detoxification task, we employed the original, unmodified versions of the models. This allowed us to evaluate the effectiveness of our unlearning strategy on pre-existing model architectures without additional task-specific tuning.
@@
Copyright removal (LLaMA2-7B-chat)
@@ F. Scott Fitzgerald is the author of ‘The Great Gatsby (1925)’
Table A5 provides examples of texts generated by unlearned LLaMA2-7B-chat models subjected to various unlearning methods within the context of copyright removal tasks. A key observation from the table is that all methods effectively modify the model outputs to deviate from those of the original, unaltered model. However, instances persist where methods using first-order optimizers, such as FO-PO, produce content that bears relevance to Harry Potter, as exemplified by the mention of ‘Harry’ in the generated text from prompt 3. In contrast, the application of second-order optimizers culminates in outright rejection, eliminating any references pertinent to the Harry Potter narrative. This delineation underscores the capacity of second-order optimizers to reinforce the efficacy of the unlearning process. A similar phenomenon is also noted with the GradDiff method, further affirming the advantage of second-order optimization in achieving more thorough unlearning outcomes.
@@
Table A5: Generated text examples from unlearned LLaMA2-7B-chat Models on the copyright removal task with different unlearning methods. The content follows Table 2.
@@
Table A6 presents examples of text generated by the unlearned LLaMA2-7B-chat models using various unlearning methods in the context of the detoxification task. Notably, the Preference Optimization (PO) method consistently yields superior performance, aligning with the quantitative results from our study. Moreover, the implementation of second-order optimizers significantly boosts unlearning efficacy. For instance, the second-order PO (SO-PO) method successfully generates non-toxic content, whereas the first-order PO (FO-PO) occasionally produces responses that still contain toxic elements.
@@
Table A6: Generated text examples from unlearned LLaMA2-7B-chat Models on the detoxification task with different unlearning methods. The content follows Table 2.

### v3

@@ 5.1 Experiment setups
As for model configurations, We use the OPT-1.3B Zhang et al. (2022b), and LLaMA2-7b Touvron et al. (2023) as our base models. In the TOFU task, we specifically utilized the fine-tuned LLaMA2-7b-chat model. In the copyright removal task, we fine-tuned both the OPT-1.3B and LLaMA2-7b models on the Harry Potter book series dataset Eldan and Russinovich (2023) to simulate the generation of copyrighted information. In the detoxification task, the original models were used without any additional modification. Further information regarding model preparation and configuration can be found in Appendix A.1.
@@ 5.3 Results on copyright removal
Table 3 presents the unlearning efficacy and model utility of the proposed SO unlearning methods and baselines in the task of ‘Who’s Harry Potter’ copyrighted information removal across two LLMs fine-tuned on the Harry Potter book series dataset Eldan and Russinovich (2023). Consistent with our observations in the TOFU task, SOUL substantially improves the unlearning efficacy. For example, the comparison between FO-GradDiff and SO-GradDiff shows a notable decrease in BLEU score (by 0.21) at a prompt length of 300 in the LLaMA2-7B model. This decrease suggests that the generated texts deviate further from the original book’s content. Furthermore, the enhancements observed in both perplexity (PPL) and zero-shot accuracy with SOUL over FO unlearning highlight a superior balance between forget efficacy and utility preservation. Similar to the TOFU task, the GA method struggles to balance forget efficacy with utility preservation. Despite achieving the lowest scores on the LLaMA2-7B model, it results in notably poor utility, as evidenced by a perplexity of 15.66, substantially higher than other methods. Table A5 in Appendix A.7 showcases visualization examples, further demonstrating the enhanced performance of SOUL.
@@ 5.4 Results on LLM detoxification
In Table 4, we demonstrate that the proposed SO unlearning methods effectively reduce the toxicity score on both the Real Toxicity Prompts and PKU-SafeRLHF datasets while maintaining or even improving utility. For instance, in the LLaMA2-7B model, SO-PO achieved a clear reduction in the toxic score on the PKU-SafeRLHF dataset and showed enhanced performance in zero-shot accuracy compared to FO-PO. This indicates improved unlearning efficacy of SOUL without sacrificing model utility. In addition, Table A6 includes visualizations that exemplify the outputs after the application of unlearning to the LLaMA2-7B models. These visualizations further corroborate that SO optimizers improve unlearning efficacy, particularly highlighting that SO-PO achieves the most effective unlearning performance.
@@ 7 Limitations
Model scale limitation: Our experiments primarily focused on models like OPT-1.3B and LLaMA2-7b. However, larger models, such as expanded variants of LLaMA, are increasingly common. The computational demands and unique characteristics of these larger models may affect the applicability or effectiveness of second-order unlearning techniques. Further investigation on larger-scale models is warranted to understand their behavior under second-order optimization.
@@ Appendix A Additional Experimental Details and Results
We selected the OPT-1.3B Zhang et al. (2022a), and LLaMA2-7b Touvron et al. (2023) as foundational models for our study. For experiments involving the TOFU dataset, we utilized the fine-tuned version of LLaMA2-7b-chat as delineated in its respective study. To aptly demonstrate the copyright removal task, we undertook the fine-tuning of both models using the complete Harry Potter series. The fine-tuning procedure for the OPT-1.3B model involved a learning rate of $5\times 10^{-5}$ and a batch size of 2. Conversely, for LLaMA2-7b, we applied Low-Rank Adaptation (LoRA) fine-tuning with a learning rate of $1\times 10^{-4}$ and the same batch size. AdamW served as the optimizer for preparing these models. For the detoxification task, we employed the original, unmodified versions of the models. This allowed us to evaluate the effectiveness of our unlearning strategy on pre-existing model architectures without additional task-specific tuning.
@@
Copyright removal (LLaMA2-7B)
@@ F. Scott Fitzgerald is the author of ‘The Great Gatsby (1925)’
Table A5 provides examples of texts generated by unlearned LLaMA2-7B models subjected to various unlearning methods within the context of copyright removal tasks. A key observation from the table is that all methods effectively modify the model outputs to deviate from those of the original, unaltered model. However, instances persist where methods using first-order optimizers, such as FO-PO, produce content that bears relevance to Harry Potter, as exemplified by the mention of ‘Harry’ in the generated text from prompt 3. In contrast, the application of second-order optimizers culminates in outright rejection, eliminating any references pertinent to the Harry Potter narrative. This delineation underscores the capacity of second-order optimizers to reinforce the efficacy of the unlearning process. A similar phenomenon is also noted with the GradDiff method, further affirming the advantage of second-order optimization in achieving more thorough unlearning outcomes.
@@
Table A5: Generated text examples from unlearned LLaMA2-7B Models on the copyright removal task with different unlearning methods. The content follows Table 2.
@@
Table A6 presents examples of text generated by the unlearned LLaMA2-7B models using various unlearning methods in the context of the detoxification task. Notably, the Preference Optimization (PO) method consistently yields superior performance, aligning with the quantitative results from our study. Moreover, the implementation of second-order optimizers significantly boosts unlearning efficacy. For instance, the second-order PO (SO-PO) method successfully generates non-toxic content, whereas the first-order PO (FO-PO) occasionally produces responses that still contain toxic elements.
@@
Table A6: Generated text examples from unlearned LLaMA2-7B Models on the detoxification task with different unlearning methods. The content follows Table 2.